OpenStudy (anonymous):

What is 3(2k-2)=-2(4k-11)

OpenStudy (anonymous):

it would be 6k-6=8k-22 i think

OpenStudy (paki):

no @mcnabkt

OpenStudy (anonymous):

-8+22*

OpenStudy (paki):

yeah now rite.

OpenStudy (paki):

now combine the like terms and solve...

OpenStudy (anonymous):

ohhh right i forgot clt!

OpenStudy (paki):

hmmm... so what will be the next step..?

OpenStudy (anonymous):

6k-2+11?

OpenStudy (paki):

6k-6=-8k+22 6k+8k= 22+6 14k= 28 k= 28/14

OpenStudy (anonymous):

ohh i was doing a different formula i think

OpenStudy (paki):

no, you were wrong in your calculations... so hope, now it will be fine...